A four digit integer, WXYZ, in which w, x, y and z each represent a different digit, is form according to the following rules.
X = W + Y + Z
W = Y + 1
Z = W - 5
WHAT IS THE FOUR-DIGIT NUMBER?

Answer by ikleyn(52873)

.

Let turn on our common sense.


Z = W - 5 must be non-negative.


It means that W is at least 5.


In turn, it means that Y is at least 4   (since W = Y+1)


And since X = W + Y + Z, it implies that


    W = 5;  Y = 4;  Z = 0;  X = 9.


ANSWER.  The number is  WXYZ = 5940.